# Approximability of Deep Computations

### v3

@@ 2.2. Deep computations and deep equilibria
Let $s_{\bullet}^{[\cdot]}$ be a confiner. Recall that $\mathcal{T}[s_{\bullet}^{[\cdot]}]$, $\mathfrak{T}[s_{\bullet}^{[\cdot]}]$ are the sets of all transforms and t-ts, respectively, that are $s_{\bullet}^{[\cdot]}$-confined. (see page 2.1 for the definitions).
@@
The CSS $\mathcal{C}$ above will often be denoted simply $\mathcal{C}=\langle\underline{L},\underline{\Gamma}\rangle$ without explicitly naming the evaluation action $\operatorname{ev}$ which, however, is always an implicit operation of $\mathcal{C}$ [^5] At any rate, the “functional application notation” $\gamma(x)$ for $\operatorname{ev}(\gamma,x)$ makes it essentially redundant to have a name for the action.
@@ 3.1. Computations, states, observable features and predicates: A meteorological allegory
Consider physical quantities (such as temperature and barometric pressure) that are real-valued, and each of which may be observed at any given point. For definiteness, consider points on or above the surface of earth, regarded as an idealized sphere. A state $v$ captures the properties a specific such point at a specific moment in time. In such idealization, each physical quantity at any $v$ is called a feature of $v$ (or observable feature for emphasis). Each such feature must be given a name (e.g., temperature, pressure, latitude, longitude, height, etc.); these names are essential, for otherwise the real value of a feature of $v$ is devoid of context. We use the term observable to refer to the name given to any such property that may be observed; in a formal treatment, we use (purely syntactic) symbols (e.g., “T” for temperature, “p” for pressure, “lat” for latitude, “long” for longitude, “h” for height, etc.) as observables. An observable feature of $v$ is the value at $v$ of the observable; e.g., $v$ may have features $\mathtt{lat}(v)=+29.42$ (the lat-feature—i.e., latitude—of $v$ is $29.42^{\circ}$ N), $\mathtt{long}(v)=-98.49$ ($v$ has long-feature—i.e., longitude—$98.49^{\circ}$ W), $\mathtt{h}(v)=229$ ($v$ is at $229$ m height) $\mathtt{T}(v)=33.5$ (the temperature at $v$ is $33.5^{\circ}$), etc.
@@
Taking $L$ together with the predicate $P(\cdot)$ interpreting each observable $P\in\mathcal{P}$ thereon, we obtain a pair $\underline{L}\coloneqq\langle L,(P(\cdot))_{\mathtt{P}\in\mathcal{P}}\rangle$ called a Computation States Structure (CSS) in 3.2.1 below. (In $\underline{L}$, the collection of predicates $P(\cdot):L\to\mathbb{R}$ is a family indexed by symbols P.) By an abuse of notation, we may denote such structure in the form $\underline{L}=\langle L,\mathcal{P}\rangle$ wherein the collection $(P(\cdot):P\in\mathcal{P})$ of predicates is implicitly identified with the indexing set $\mathcal{P}$.
In the allegory, such features include the quantities lat$(v)$, long$(v)$ and h$(v)$, which are coordinates in the usual sense, as well as other features T$(v)$, p$(v)$ and time t$(v)$, which are not; however, this suggests regarding the collection of all features $(P(v))_{P\in\mathcal{P}}$ of states as coordinatizing states $v\in L$. Each $P$-feature $P(v)$ is the “$P$-th coordinate” of $v$ in an abstract sense; the collection $\operatorname{tp}(v)\coloneqq(P(v))_{P\in\mathcal{P}}$ is called the type of $v$. Any state $v$ is uniquely characterized by its type. A critical feature of our approach is to endow the state space $L$ with the topology of “pointwise convergence”, i.e., a filter on (or: a sequence or net of) states converges to a state $v\in L$ iff the filter (or sequence, or net) of real-valued $P$-features converges to $P(v)$, for each $P\in\mathcal{P}$.
@@
For illustration purposes, consider the “advance-time-by-1” computation $\alpha$ taking any state $v$ of some point at some time $(v)=t$ to the unique state $w=\alpha(v)$ of the same point at time t$(w)=t+1$. Features $T(\alpha),p(\alpha)$ of the computation $\alpha$ give the temperature and pressure at a future moment $t+1$ in time from the state at present time $t$. Meteorologists would be ecstatic to learn features at time $t+1$ from those at time $t$!
@@
where $\underline{L}$ is a CSS, and $\underline{\Gamma}=\langle\Gamma,\circ\rangle$ is any semigroup under an (associative) composition operation $\circ$, with elements $\gamma\in\Gamma$ representing computations on $L$ via an evaluation map $\operatorname{ev}:\Gamma\times L\to L:(\gamma,v)\mapsto\operatorname{ev}(\gamma,v)$ ($=\gamma(v)$, if $\Gamma$ is already a set of maps $\gamma:L\to L$). Layer state transitions $\gamma(\cdot):v\mapsto\operatorname{ev}(\gamma,v)$ are assumed continuous on $L$ (when $L$ is noncompact, we require them to be extendable in the sense of §4.2.1). CCSs are the natural structures to study compositions $\gamma_{n}\circ\gamma_{n-1}\circ\dots\circ\gamma_{2}\circ\gamma_{1}$ of $n$-many computations leading, as $n\to\infty$, to “deep computation states”, as well as “deep iterates” asymptotically approximated by $n$-fold iterates $\gamma^{(n)}=\gamma\circ\gamma\circ\dots\circ\gamma$ of a fixed computation $\gamma$.
@@ A Computation States Structure (CSS) with predicates $\mathcal{P}$ is of the form
  • For each symbol $P\in\mathcal{P}$, the $P$-predicate of $\underline{L}$ is a real function $P(\cdot):L\to\mathbb{R}$ [^10] In the setting of real-valued structures, any real function is called a predicate.
@@ 3.2.1. Types of states
Ultrafilters on an infinite (“index”) set $I$ will be denoted $\mathcal{U},\mathcal{V},\dots$; we consider nonprincipal ultrafilters tacitly. Given an ultrafilter $\mathcal{U}$ on $I$, we say that an indexed family $\mathfrak{v}^{(\bullet)}\coloneqq(\mathfrak{v}^{(i)})_{i\in I}$ of elements $\mathfrak{v}^{(i)}=(\mathfrak{v}^{(i)}_{P})_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ converges to $\mathfrak{u}=(\mathfrak{u}_{P})_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$, or that $\mathfrak{u}$ is the $\mathcal{U}$-ultralimit of $\mathfrak{v}^{(\bullet)}$ with respect to $\mathcal{U}$ if $\operatorname{\mathcal{U}lim}_{i}\mathfrak{v}_{P}^{(i)}=\mathfrak{u}_{P}$ for each $P\in\mathcal{P}$ (i.e., when $\mathfrak{u}$ is the $\mathcal{U}$-ultralimit of $(\mathfrak{v}^{(i)})$ in the pointwise convergence topology—not necessarily uniformly as $P$ varies). [^11] When it exists, the $\mathcal{U}$-ultralimit $s=\operatorname{\mathcal{U}lim}_{i}r_{i}$ of $(r_{i})_{i\in I}\subseteq\mathbb{R}$ is uniquely characterized by the following property: for every $\varepsilon>0$, the set $\{i\in I:\left|r_{i}-s\right|<\varepsilon\}$ belongs to $\mathcal{U}$ (i.e., is a “$\mathcal{U}$-large” set). Not all ultralimits $\operatorname{\mathcal{U}lim}_{i}P(u^{{(i)}})$ need exist since $\mathbb{R}$ is not compact. The (necessarily unique) ultralimit $\mathfrak{u}$ is denoted $\mathcal{U}\lim_{i}\mathfrak{v}^{(i)}$.
@@ 3.2.2. Topology on the layer state space
Even if not imposed a priori on a CSS $\underline{L}$, the Reduction Axiom is always satisfied if one replaces $L$ by its quotient $\tilde{L}\coloneqq L/{\operatorname{tp}}$ upon identifying equal-in-type states, and each predicate $P(\cdot)$ on $L$ by the naturally induced predicate $\tilde{P}(\cdot)$ on $\tilde{L}$. From a structural viewpoint, $\langle L,(P(\cdot))_{P\in\mathcal{P}}\rangle$ and $\langle\tilde{L},(\tilde{P}(\cdot))_{P\in\mathcal{P}}\rangle$ are identical (isomorphic, in the sense of Keisler’s General Real-Valued Structures [Kei23]).
@@ 3.3.1. Tychonoff spaces
The distinguished predicates $P(\cdot)$ are regarded as being “computable on $L$ab initio; they also may be seen as monomials generating some polynomial algebra of continuous real functions on $L$; the uniform closure of the set of such monomials is the algebra $\mathbf{D}$ of “definable predicates” on $L$ (which are, by necessity, continuous real functions on $L$). In general, however, $\mathbf{D}$ is a proper subalgebra of the full algebra $\mathrm{C}(L)\supseteq\mathbf{D}$ of continuous real functions on $L$. Any function $\varphi\in\mathrm{C}(L)\setminus\mathbf{D}$ is non-definable over $\mathcal{P}$; it is appropriate to think of such $\varphi$ as “transcendental” over $\mathcal{P}$ —not merely in an algebraic sense, but in a stronger topological one: not only does such a non-definable $\varphi$ fail to be a polynomial on monomials $P(\cdot)\in\mathcal{P}$; in fact, it is not even possible to approximate $\varphi$ uniformly on $L$ by such polynomials.
@@ 3.3.3. Realcompactness of type spaces
The type space $\mathfrak{X}=\mathfrak{X}_{\Phi}$ depends on $\Phi$. A key observation is that each of the functions $\varphi\in\Phi$ extends to $\mathfrak{X}$ continuously (as the “$\varphi$-th coordinate” map $\hat{\varphi}:(\mathfrak{x}_{\psi})_{\psi\in\Phi}\mapsto\mathfrak{x}_{\varphi}$). However, other real functions on $X$—even if continuous—need not extend to $\mathfrak{X}$ continuously. Thus, the $\Phi$-type space $\mathfrak{X}\supseteq\Phi\text{-}\!\operatorname{tp}(X)\cong X$ possesses the universal property that every $\varphi\in\Phi$ admits a unique continuous extension to $\mathfrak{X}$; it is characterized by such universal property up to homeomorphism. [^12] In fact, any real function $\xi$ on $X$ that is uniformly approximable by polynomials in the functions $\varphi\in\Phi$ is (necessarily continuous, and) extends continuously to a real function $\hat{\xi}$ on $\mathfrak{X}$ (a uniform limit of polynomials in the corresponding functions $\hat{\varphi}$ on $\mathfrak{X}$), so $\mathfrak{X}$ possesses the extension property for functions in the uniform closure of the real algebra generated by functions $\varphi\in\Phi$.
@@
Let $C\coloneqq\mathrm{C}(X)$ be the set of all continuous real-valued functions on a Tychonoff space $X$, and let $\hat{X}\subseteq\mathbb{R}^{C}$ be the corresponding type space, called the realcompactification of $X$. Every continuous function $\varphi$ on $X$ extends to a continuous function $\hat{\varphi}$ on the realcompactification $\hat{X}$. In fact, for any $\Phi\subseteq C=\mathrm{C}(X)$, the $\Phi$-type space $\mathfrak{X}\subseteq\mathbb{R}^{\Phi}$ is a quotient (not a subspace!) of $\hat{X}\subseteq\mathbb{R}^{C}$ in a natural manner: indeed, $\mathfrak{X}\subseteq\mathbb{R}^{\Phi}$ is the image of $\hat{X}\subseteq\mathbb{R}^{C}$ under the natural projection map $\mathbb{R}^{C}\to\mathbb{R}^{\Phi}:(\mathfrak{x}_{\varphi}:\varphi\in C)\mapsto(\mathfrak{x}_{\varphi}:\varphi\in\Phi)$. Thus, given a fixed set $\Phi$ of continuous real functions on $X$, it is appropriate to think of the $\Phi$-type space $\mathfrak{X}$ as a “$\Phi$-relative realcompactification” of $X$, since $\mathfrak{X}\supseteq X$ possesses the universal extension property only for functions $\varphi\in\Phi$ —rather than for all $\varphi\in\mathrm{C}(X)$, which corresponds to the (“absolute”) realcompactification $\hat{X}$ of $X$ [^13] The notation $\upsilon X$ (“upsilon-$X$”) is standard for the realcompactification—denoted $\hat{X}$ above—of a Tychonoff space $X$.
@@ 3.3.4. Realcompact CSSs
It is appropriate to regard realcompactness as capturing a certain notion of “completeness” or “saturation” of the space $L$. Particularly when $\mathcal{P}$ is infinite, realcompactness is a rather strong requirement on CCSs, so we do not impose it as an axiom; instead, we rely primarily on the realcompactness and universal properties of the type space $\mathfrak{L}\supseteq L$ [^14] When $\mathcal{P}$ is finite, realcompactness is a rather mild requirement: it is seen to be equivalent to the completeness of $\langle L,\delta\rangle$, where $\delta$ is the metric introduced in the proof of Proposition 2.1.
@@ A Compositional Computation Structure (CCS)
  • a semigroup Γ¯=($\underline{\Gamma}=(\Gamma,\circ)$, the computations sort (the—associative—semigroup operation $\circ:\Gamma\times\Gamma\to\Gamma$ is denoted simply $(\gamma,\delta)\mapsto\gamma\delta$ when convenient);
@@
    In principle, the semigroup operation $\circ$’ of $\underline{\Gamma}$ and evaluation action ‘$\operatorname{ev}$’ are abstract (i.e., not literally composition and application of functions). However, one may always regard $\underline{\Gamma}$ as a semigroup (under the operation $\circ$’ interpreted as composition) of maps $\gamma(\cdot):v\mapsto\operatorname{ev}(\gamma,v)$; —i.e., regard $\Gamma$ as a sub-semigroup of the semigroup $L^{L}$ of all maps $L\to L$, under bona fide functional composition: Nothing of structural relevance is lost thus. The structural viewpoint abstracts inessential aspects of a concrete such realization of $\Gamma$. In practice, it is convenient to identify $\gamma\in\Gamma$ with $\gamma(\cdot)\in L^{L}$.

### v5

@@ 1.2. Deep computations and deep equilibria
Let $s_{\bullet}^{[\cdot]}$ be a confiner. Recall that $\mathcal{T}[s_{\bullet}^{[\cdot]}]$, $\mathfrak{T}[s_{\bullet}^{[\cdot]}]$ are the sets of all transforms and t-ts, respectively, that are $s_{\bullet}^{[\cdot]}$-confined. (see page 1.1 for the definitions).
@@
The CSS $\mathcal{C}$ above will often be denoted simply $\mathcal{C}=\langle\underline{L},\underline{\Gamma}\rangle$ without explicitly naming the evaluation action $\operatorname{ev}$ which, however, is always an implicit operation of $\mathcal{C}$. [^3] At any rate, the “functional application notation” $\gamma(x)$ for $\operatorname{ev}(\gamma,x)$ makes it essentially redundant to have a name for the action.
@@ 2.1. Computations, states, observable features and predicates: A meteorological allegory
[…]
@@
Taking $L$ together with the predicate $P(\cdot)$ interpreting each observable $P\in\mathcal{P}$ thereon, we obtain a pair $\underline{L}\coloneqq\langle L,(P(\cdot))_{\mathtt{P}\in\mathcal{P}}\rangle$ called a Computation States Structure (CSS) in 2.2.1 below. (In $\underline{L}$, the collection of predicates $P(\cdot):L\to\mathbb{R}$ is a family indexed by symbols P.) By an abuse of notation, we may denote such structure in the form $\underline{L}=\langle L,\mathcal{P}\rangle$ wherein the collection $(P(\cdot):P\in\mathcal{P})$ of predicates is implicitly identified with the indexing set $\mathcal{P}$.
In the allegory, such features include the quantities lat(v)(v), long(v)(v) and h(v)(v), which are coordinates in the usual sense, as well as other features T(v)(v), p(v)(v) and time t(v)(v), which are not; however, this suggests regarding the collection of all features $(P(v))_{P\in\mathcal{P}}$ of states as coordinatizing states $v\in L$. Each $P$-feature $P(v)$ is the “$P$-th coordinate” of $v$ in an abstract sense; the collection $\operatorname{tp}(v)\coloneqq(P(v))_{P\in\mathcal{P}}$ is called the type of $v$. Any state $v$ is uniquely characterized by its type. A critical feature of our approach is to endow the state space $L$ with the topology of “pointwise convergence”, i.e., a filter on (or: a sequence or net of) states converges to a state $v\in L$ iff the filter (or sequence, or net) of real-valued $P$-features converges to $P(v)$, for each $P\in\mathcal{P}$.
@@
For illustration purposes, consider the “advance-time-by-1” computation $\alpha$ taking any state $v$ of some point at some time t(v)=t(v)=t to the unique state $w=\alpha(v)$ of the same point at time t(w)=t+1(w)=t+1. Features $T(\alpha),p(\alpha)$ of the computation $\alpha$ give the temperature and pressure at a future moment $t+1$ in time from the state at present time $t$. Meteorologists would be ecstatic to learn features at time $t+1$ from those at time tt!
@@
where $\underline{L}$ is a CSS, and $\underline{\Gamma}=\langle\Gamma,\circ\rangle$ is any semigroup under an (associative) composition operation $\circ$, with elements $\gamma\in\Gamma$ representing computations on $L$ via an evaluation map $\operatorname{ev}:\Gamma\times L\to L:(\gamma,v)\mapsto\operatorname{ev}(\gamma,v)$ ($\gamma(v)$, if $\Gamma$ is already a set of maps $\gamma:L\to L$). Layer state transitions $\gamma(\cdot):v\mapsto\operatorname{ev}(\gamma,v)$ are assumed continuous on $L$ (when $L$ is noncompact, we require them to be extendable in the sense of §3.2.1). CCSs are the natural structures to study compositions $\gamma_{n}\circ\gamma_{n-1}\circ\dots\circ\gamma_{2}\circ\gamma_{1}$ of $n$-many computations leading, as $n\to\infty$, to “deep computation states”, as well as “deep iterates” asymptotically approximated by $n$-fold iterates $\gamma^{(n)}=\gamma\circ\gamma\circ\dots\circ\gamma$ of a fixed computation $\gamma$.
@@ A Computation States Structure (CSS) with predicates $\mathcal{P}$ is of the form
  • For each symbol $P\in\mathcal{P}$, the $P$-predicate of $\underline{L}$ is a real function $P(\cdot):L\to\mathbb{R}$. [^8] In the setting of real-valued structures, any real function is called a predicate.
@@ 2.2.1. Types of states
Ultrafilters on an infinite (“index”) set $I$ will be denoted $\mathcal{U},\mathcal{V},\dots$; we consider nonprincipal ultrafilters tacitly. Given an ultrafilter $\mathcal{U}$ on $I$, we say that an indexed family $\mathfrak{v}^{(\bullet)}\coloneqq(\mathfrak{v}^{(i)})_{i\in I}$ of elements $\mathfrak{v}^{(i)}=(\mathfrak{v}^{(i)}_{P})_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ converges to $\mathfrak{u}=(\mathfrak{u}_{P})_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$, or that $\mathfrak{u}$ is the $\mathcal{U}$-ultralimit of $\mathfrak{v}^{(\bullet)}$ with respect to $\mathcal{U}$ if $\operatorname{\mathcal{U}lim}_{i}\mathfrak{v}_{P}^{(i)}=\mathfrak{u}_{P}$ for each $P\in\mathcal{P}$ (i.e., when $\mathfrak{u}$ is the $\mathcal{U}$-ultralimit of $(\mathfrak{v}^{(i)})$ in the pointwise convergence topology—not necessarily uniformly as $P$ varies). [^9] When it exists, the $\mathcal{U}$-ultralimit $s=\operatorname{\mathcal{U}lim}_{i}r_{i}$ of $(r_{i})_{i\in I}\subseteq\mathbb{R}$ is uniquely characterized by the following property: for every $\varepsilon>0$, the set $\{i\in I:\left|r_{i}-s\right|<\varepsilon\}$ belongs to $\mathcal{U}$ (i.e., is a $\mathcal{U}$-large” set). Not all ultralimits $\operatorname{\mathcal{U}lim}_{i}P(u^{{(i)}})$ need exist since $\mathbb{R}$ is not compact. The (necessarily unique) ultralimit $\mathfrak{u}$ is denoted $\mathcal{U}\lim_{i}\mathfrak{v}^{(i)}$.
@@ 2.2.2. Topology on the layer state space
[…]
@@ 2.3.1. Tychonoff spaces
The distinguished predicates $P(\cdot)$ are regarded as being “computable on LLab initio; they also may be seen as monomials generating some polynomial algebra of continuous real functions on $L$; the uniform closure of the set of such monomials is the algebra $\mathbf{D}$ of “definable predicates” on $L$ (which are, by necessity, continuous real functions on $L$). In general, however, $\mathbf{D}$ is a proper subalgebra of the full algebra $\mathrm{C}(L)\supseteq\mathbf{D}$ of continuous real functions on $L$. Any function $\varphi\in\mathrm{C}(L)\setminus\mathbf{D}$ is non-definable over $\mathcal{P}$; it is appropriate to think of such $\varphi$ as “transcendental” over $\mathcal{P}$ —not merely in an algebraic sense, but in a stronger topological one: not only does such a non-definable $\varphi$ fail to be a polynomial on monomials $P(\cdot)\in\mathcal{P}$; in fact, it is not even possible to approximate $\varphi$ uniformly on $L$ by such polynomials.
@@ 2.3.3. Realcompactness of type spaces
The type space $\mathfrak{X}=\mathfrak{X}_{\Phi}$ depends on $\Phi$. A key observation is that each of the functions $\varphi\in\Phi$ extends to $\mathfrak{X}$ continuously (as the $\varphi$-th coordinate” map $\hat{\varphi}:(\mathfrak{x}_{\psi})_{\psi\in\Phi}\mapsto\mathfrak{x}_{\varphi}$). However, other real functions on XX—even if continuous—need not extend to $\mathfrak{X}$ continuously. Thus, the $\Phi$-type space $\mathfrak{X}\supseteq\Phi\text{-}\!\operatorname{tp}(X)\cong X$ possesses the universal property that every $\varphi\in\Phi$ admits a unique continuous extension to $\mathfrak{X}$; it is characterized by such universal property up to homeomorphism. [^10] In fact, any real function $\xi$ on $X$ that is uniformly approximable by polynomials in the functions $\varphi\in\Phi$ is (necessarily continuous, and) extends continuously to a real function $\hat{\xi}$ on $\mathfrak{X}$ (a uniform limit of polynomials in the corresponding functions $\hat{\varphi}$ on $\mathfrak{X}$), so $\mathfrak{X}$ possesses the extension property for functions in the uniform closure of the real algebra generated by functions $\varphi\in\Phi$.
@@
Let $C\coloneqq\mathrm{C}(X)$ be the set of all continuous real-valued functions on a Tychonoff space $X$, and let $\hat{X}\subseteq\mathbb{R}^{C}$ be the corresponding type space, called the realcompactification of $X$. Every continuous function $\varphi$ on $X$ extends to a continuous function $\hat{\varphi}$ on the realcompactification $\hat{X}$. In fact, for any $\Phi\subseteq C=\mathrm{C}(X)$, the $\Phi$-type space $\mathfrak{X}\subseteq\mathbb{R}^{\Phi}$ is a quotient (not a subspace!) of $\hat{X}\subseteq\mathbb{R}^{C}$ in a natural manner: indeed, $\mathfrak{X}\subseteq\mathbb{R}^{\Phi}$ is the image of $\hat{X}\subseteq\mathbb{R}^{C}$ under the natural projection map $\mathbb{R}^{C}\to\mathbb{R}^{\Phi}:(\mathfrak{x}_{\varphi}:\varphi\in C)\mapsto(\mathfrak{x}_{\varphi}:\varphi\in\Phi)$. Thus, given a fixed set $\Phi$ of continuous real functions on $X$, it is appropriate to think of the $\Phi$-type space $\mathfrak{X}$ as a $\Phi$-relative realcompactification” of $X$, since $\mathfrak{X}\supseteq X$ possesses the universal extension property only for functions $\varphi\in\Phi$ —rather than for all $\varphi\in\mathrm{C}(X)$, which corresponds to the (“absolute”) realcompactification $\hat{X}$ of XX. [^11] The notation $\upsilon X$ (“upsilon-XX”) is standard for the realcompactification—denoted $\hat{X}$ above—of a Tychonoff space $X$.
@@ 2.3.4. Realcompact CSSs
It is appropriate to regard realcompactness as capturing a certain notion of “completeness” or “saturation” of the space $L$. Particularly when $\mathcal{P}$ is infinite, realcompactness is a rather strong requirement on CCSs, so we do not impose it as an axiom; instead, we rely primarily on the realcompactness and universal properties of the type space $\mathfrak{L}\supseteq L$. [^12] When $\mathcal{P}$ is finite, realcompactness is a rather mild requirement: it is seen to be equivalent to the completeness of $\langle L,\delta\rangle$, where $\delta$ is the metric introduced in the proof of Proposition 1.1.
@@ A Compositional Computation Structure (CCS)
  • a semigroup $\underline{\Gamma}=(\Gamma,\circ)$, the computations sort (the—associative—semigroup operation $\circ:\Gamma\times\Gamma\to\Gamma$ is denoted simply $(\gamma,\delta)\mapsto\gamma\delta$ when convenient);
@@
    In principle, the semigroup operation $\circ$’ of $\underline{\Gamma}$ and evaluation action $\operatorname{ev}$’ are abstract (i.e., not literally composition and application of functions). However, one may always regard $\underline{\Gamma}$ as a semigroup (under the operation $\circ$’ interpreted as composition) of maps $\gamma(\cdot):v\mapsto\operatorname{ev}(\gamma,v)$; —i.e., regard $\Gamma$ as a sub-semigroup of the semigroup $L^{L}$ of all maps $L\to L$, under bona fide functional composition: Nothing of structural relevance is lost thus. The structural viewpoint abstracts inessential aspects of a concrete such realization of $\Gamma$. In practice, it is convenient to identify $\gamma\in\Gamma$ with $\gamma(\cdot)\in L^{L}$.